OpenStudy (anonymous):

Miguel is designing shipping boxes that are rectangular prisms. the shape of one box with height h in feet has a volume defined by the function V(h)=h(h-10)(h-8) what is the domain 0

OpenStudy (anonymous):

V(h) = h(h - 10)(h - 8) V(h) = h^3 - 18h^2 + 80h 0 = 3h^2 - 36h + 80 when you use the quadratic formula you get h = 2.9449495367 --or-- h = 9.0550504633

OpenStudy (anonymous):

V''(h) = 6h - 36

OpenStudy (anonymous):

V''(2.9449495367) = -18.33 V''(9.0550504633) = 18.33

OpenStudy (anonymous):

the maximum volume occurs at h=2.9449495367. Stick that into V(h) and you get: V(2.9449495367) = 105.0276 your answer would be 105ft^3

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

let me know your score ( :

OpenStudy (anonymous):

thank you sooo much!

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

You're welcome !

OpenStudy (anonymous):

if i need help again can i ask you? if you are online.

OpenStudy (anonymous):

Yep, I'll be on tomorrow. what time is it where you live ?

OpenStudy (anonymous):

10:23 pm

OpenStudy (anonymous):

Lol true. It's 11:23pm here

OpenStudy (anonymous):

lol thanks again!

OpenStudy (anonymous):

Yep. if you need help tomorrow just @ me

OpenStudy (anonymous):

i might be on tomorrow not sure yet! your awesome! alright good night!

OpenStudy (anonymous):

Night

OpenStudy (tiredbutradiantt):

TYSM ^^